determine whether the sequence is arithmetic if it is find the common difference 3,5/2,2,3/2,1,1/2

Answer by stanbon(75887)

determine whether the sequence is arithmetic if it is find the common difference 3,5/2,2,3/2,1,1/2
-------
Subtract the pairs as they progress to see if there is a common difference:
(5/2)-3 = -1/2
----
2-(5/2) = -1/2
----
(3/2)- 2 = -1/2
----
1-(3/2) = -1/2
----
(1/2)-1 = -1/2
---
It is arithmetic and the common difference is -1/2
=======================
Cheers,
Stan H.
=========================